[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Jun 22, 2016
nhc.tsbadmin at noaa.gov
Wed Jun 22 13:19:20 UTC 2016
Here is today's summary of NHC computer operations:
The situational monitor issue was resolved by rebooting the PC that
controls it and restarting the applications in the four monitors. Kim will
make restarting the applications part of the weekly update procedure to
ensure the updates worked properly.
The spurious yellow line on one of the ATCF systems was an intermittent
problem with the monitor itself, rather than the software. If this becomes
more of a problem, TSB will re-address this issue.
TSB is still working with NCO to restore the sounding processing and other
problems due to the disk failure on Mike's Linux workstation.
Matt opened an NCF ticket to find the source of the spurious pre-TCM and
TCM messages. I
Salim Leyva will be at NHC on Thursday along with an NCO contractor to help
restore the local GOES-east ingest system. They will also assist with
installing the new CARCAH UPS system.
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Ncep.nhc.nco_contacts